Find the general solution of the equation d2y/dx2+4dy/dx+4y=4e-2x

1
Expert's answer
2021-08-04T18:20:45-0400

Related homogeneous differential equation


"y''+4y'+4y=0"

The roots of the characteristic equation are


"r^2+4r+4=0"

"(r+2)^2=0"

"r_1=r_2=-2"

The general solution of the homogeneous differential equation is


"y_h=c_1e^{-2x}+c_2xe^{-2x}"

Find the particular solution of the nonhomogeneous differential equation


"y_p=Ax^2 e^{-2x}"

"y_p'=2Axe^{-2x}-2Ax^2e^{-2x}"

"y_p''=2Ae^{-2x}-8Axe^{-2x}+4Ax^2e^{-2x}"

Substitute


"2Ae^{-2x}-8Axe^{-2x}+4Ax^2e^{-2x}"

"+8Axe^{-2x}-8Ax^2e^{-2x}+4Ax^2 e^{-2x}=4e^{-2x}"


"A=2"

"y_p=2x^2 e^{-2x}"

The general solution of the nonhomogeneous differential equation is


"y=y_h+y_p"

"y=c_1e^{-2x}+c_2xe^{-2x}+2x^2 e^{-2x}"